OpenStudy (bibby):

\(3^-1 mod 59\) how do negative exponents and modular arithmetic work?

jimthompson5910 (jim_thompson5910):

3^-1 is the multiplicative inverse of 3 in mod 59 so you are looking for an x value that makes this equation true 3x = 1 (mod 59)

OpenStudy (bibby):

ah, so 3^-5 would be 3^5x?

OpenStudy (bibby):

what does it mean when it says \(3^{-5}\cong 20\)

OpenStudy (bibby):

mod 59, but that isn't important (I don't think)

ganeshie8 (ganeshie8):

we know that \(3^{-1}\) is \(\dfrac{1}{3}\) in modulo \(59\) : \[\dfrac{1}{3} = \dfrac{20}{60}\equiv \dfrac{20}{1} = 20\]

OpenStudy (bibby):

oh yeah, you showed me this trick a week ago

OpenStudy (bibby):

and so \(3^{-5}=(3^{-1})^5\)

OpenStudy (bibby):

which is congruent to 17, which is what my professor got. awesome. thank you so much ganesh. sorry about being a jerk jim

OpenStudy (bibby):

and the other 2